Allan walked from his home two thirds of the total distance to town and stole a ride for three fourths of the remaining distance. He then had one-tenth of a km left to go. How far from the town did Allan live?
Answer by CubeyThePenguin(3113)

2x/3 + x/4 + 10 = x

11x/12 + 10 = 12x/12

x/12 = 10

x = 120 km